# In Search of the Long-Tail: Systematic Generation of Long-Tail Inferential Knowledge via Logical Rule Guided Search

### v2

@@ 3.1 Generating Long-tail Statements from Symbolic Rules
We generate long-tail statements from a subset of 200 symbolic rules from LINT using ChatGPT and GPT4. We prompt the LLMs with a natural language version of the symbolic rule with data types of the variables specified, and ask them to populate the rule with all variables simultaneously. Table 2 shows the corresponding prompt of the symbolic rule example in § 2.2. To ensure long-tailness, we add “Use less frequent terms of A and B and C” in the prompt.  [^7] Other prompts such as “Use terms that are less common” and “Use terms that have lower probability in language model distribution” have similar effect. For each rule, we obtain 200 statements in each of head and long-tail distribution, from ChatGPT and GPT4 respectively.
3.2 LLMs Cannot Reach the Long-tail Distribution with Only Instructions
We calculate the log likelihood over InstructGPT of all statements generated by LINK, ChatGPT and GPT4. We calculate $\delta=mean(D(H))-mean(D(L))$ for each set of statements generated from each symbolic rule, where $D(\cdot)$ means the log likelihood distribution of InstructGPT, $H$ is the set of statements intended to be the head distribution, and $L$ is the set of the statements intended to be in the long-tail distribution.
@@ 3.2 LLMs Cannot Reach the Long-tail Distribution with Only Instructions
Figure 3 illustrates how ChatGPT and GPT4 are not able to generate long-tail statements merely from prompting. Each grid on the x-axis represents a unique symbolic rule, and each grid on the y-axis represents $\delta\in[-1.5,2]$. A $\delta$ close to 0 means that the intended head and long-tail distribution are inseparable, being larger than 0.3 empirically means a decent separation, and being negative means the intended head and long-tail distributions are reverted on the log likelihood scale of InstructGPT.
Averaged across 200 sampled rules, LINT has a positive $\delta$ of 0.48, while ChatGPT and GPT4 each has a delta of -0.14 and -0.02. The $\delta$ values for LINT (blue line) float above 0 for most of the rules, some even being above 0.5. On the other hand, $\delta$ values for ChatGPT (red line) and GPT4 (green line) mostly locate around 0, with many being negative. Low and negative $\delta$ values indicate that LLMs’ generations are in fact not in the long-tail distribution, and sometimes even in the head distribution.
To better illustrate the distribution of statements generated by ChatGPT and GPT4, we plot the log likelihood of the generated statements for one symbolic rule from the three methods Figure 4. To eliminate noise from incorrect statements on the distribution, we only plot the statements that are marked as correct in human evaluation (explained in § 3.3). The likelihood distribution of more rules can be found in Figure 9.
The long-tail statements from LINK clearly fall in a much lower probability distribution than GPT4’s generations. Moreover, GPT4’s generation in the “long-tail distribution” in fact falls in the same distribution as its head distribution.
Ablation on the reranker model.
Rerankers are essential for pushing the generations into the long-tail distribution. Figure 6 presents the distribution comparison of generated statements by LINK and the variant without the reranker. Without the reranker, the generated statements for both head distribution and long-tail distribution are pulled towards the center of the distribution, making them completely inseparable.
Post-hoc reranking over the LLM generated knowledge does not have the same effect as step-wise reranking in LINK. Figure 7 illustrates the distribution of generated statements by LINK, compared to prompt-based GPT4 and prompt-based GPT4 reranked by InstructGPT. Post-hoc reranking barely changed the distribution of generations, even with the same model as the evaluation.
Even though log likelihood is both used by reranker and the evaluation, they are taken from different models and using different inputs. Figure 10 shows that although statements generally tend to fall in the same distribution of llama-7B and InstructGPT, not all rules have cleanly separable generations. In addition, the statements we use for ranking the knowledge beams are shorter than the final statement, as they only consist of partial predicates. Despite these differences, step-wise reranking using a smaller model achieves the separation that post-hoc filtering with the evaluation model cannot achieve.
Our findings highlight the importance of performing variable-level reranking in LINK.
3.3 LLMs Struggle with Data Correctness with Only Instructions
@@
In addition to distribution correctness, we also evaluate data type conformity and factual correctness of LLMs’ long-tail knowledge generations using crowdworkers from Amazon Mechanic Turk [^8] https://www.mturk.com (AMT). For data type conformity, we ask an AMT worker Is {variable} a {data type}? for each variable in the symbolic rule. For factual correctness, we ask an AMT worker Does the premise entail the conclusion? We sample 4,000 statements from LINT for human evaluation, of which 2,025 are from head distribution and 1,975 are from long-tail distribution. We take 3 annotations for each statement and take the majority vote. Annotator agreement can be found in Appendix G.3. The AMT template can be found in Appendix G.2.
Table 3 shows that prompt-based ChatGPT and GPT4 underperform LINK in both data type conformity and factual correctness. Without LINK, both models struggle more with factual correctness, a foreseeable behavior in the low likelihood realm. For domain wise performance see Table 14. For examples of failure cases, see Table 17.
@@
Critic models are essential for guaranteeing the generation quality, especially in the long-tail distribution. Table 8 in Appendix C.2 shows that removing the critic and removing both reranker and critic leads to significant drops of data type conformity and factual correctness of generations in the long-tail distribution. Note that LINK w/o reranker + critic has higher generation quality than LINK w/o critic because the former framework cannot generate long-tail statements without the reranker, so the generations are in the head distribution. This observation further suggests that critic models are essential for generation qualities in the long-tail distribution.
3.4 LINK Facilitates Long-tail Generation Regardless of the Knowledge Model
Our analysis above has shown that by simply prompting ChatGPT and GPT4, arguably the state-of-the-art LLMs on language tasks, we cannot effectively generate knowledge statements that are both high quality and in the long-tail distribution. By adding LINK, we can improve both distribution correctness and generation quality, as shown in Figure 8 and Table 10.
@@ 3.4 LINK Facilitates Long-tail Generation Regardless of the Knowledge Model
Domain Distribution Template InstructGPT ChatGPT GPT4 Human Baseline
Zero-Shot COT Zero-Shot COT Zero-Shot COT
Natural Properties Head Positive Template 3.47 18.52 33.54 35.95 31.33 50.13 -
Negative Template 93.85 70.75 72.68 60.96 89.05 81.58 -
All 52.13 46.64 54.62 49.41 62.41 67.07 82.31
Longtail Positive Template 3.71 16.97 33.39 34.09 24.62 44.53 -
Negative Template 93.84 70.50 71.21 60.13 87.36 79.57 -
All 52.24 45.79 53.75 48.11 58.40 63.40 82.45
Temporal Head Positive Template 12.12 47.96 45.28 59.83 79.43 87.03 -
Negative Template 88.23 79.33 83.57 80.35 86.20 79.81 -
All 53.10 64.85 65.90 70.88 83.07 83.14 84.69
Longtail Positive Template 16.56 47.01 44.29 58.72 77.41 81.99 -
Negative Template 83.89 76.02 79.90 76.92 83.08 76.39 -
All 52.81 62.63 63.46 68.52 80.46 78.97 83.20
Capability and Advice Head Positive Template 22.61 52.85 61.84 70.38 74.64 78.46 -
Negative Template 84.80 71.65 77.66 74.91 86.93 84.63 -
All 56.10 62.97 70.36 72.82 81.26 81.78 83.83
Longtail Positive Template 27.66 55.38 67.01 68.77 72.87 80.89 -
Negative Template 78.22 69.71 74.19 74.24 82.87 79.60 -
All 54.88 63.09 70.88 71.71 78.25 80.19 85.13
Locational Head Positive Template 22.86 52.86 51.67 63.10 52.14 59.29 -
Negative Template 87.55 65.92 68.98 61.84 75.71 77.14 -
All 57.69 59.89 60.99 62.42 64.84 68.90 75.71
Longtail Positive Template 41.39 57.30 51.87 57.68 59.74 63.30 -
Negative Template 76.24 62.12 63.40 58.11 73.03 70.30 -
All 60.16 59.90 58.08 57.91 66.90 67.07 67.42
@@
Table 4: Performance on the entailment classification task of three LLMs decreases on the long-tail distribution compared to the head distribution, while human performance does not. LLMs also have a template bias towards negative templates – when the target token is “no”, “false” and “wrong”.
4 Can LLM Reason Effectively in the Long-tail Distribution?
@@
We used all human evaluated knowledge statements that are labeled correct or factually incorrect in LINT, with 1,925 statements in head distribution and 1,856 statements in long-tail distribution, and created a simple entailment classification task to motivate usage of our LINT dataset.
We evaluate three LLMs: InstructGPT, ChatGPT and GPT4, using both zero-shot and Chain-of-Thought (COT) (Wei et al., 2022) prompting.
We use data labeled as both factually correct and conforming data type as positive samples and use data labeled as factually incorrect during human evaluation as negative samples. For all positive samples, entailment forms between the premise and conclusion, while for negative samples, non-entailment forms. Given that LINT has an 84% accuracy, we have more positive samples than negative samples.
@@
In order to examine whether data distribution affects reasoning behavior in LLMs, we convert each statement into 13 probing questions, where each probing template corresponds to either a positive label (“Yes”, “True”, or “Right”) or a negative label (“No”, “False”, or “Wrong”). The probing templates are summarized in Appendix E.1.
@@
We obtain human performance on the same set of statements. We recruit 17 AMT workers who do not participate in the evaluation task (and thus have not seen the task data). The workers see the knowledge statements in premise, conclusion format and are asked to select “yes/no” to whether the premise entails the conclusion. The workers are asked to use search engines to verify their answers. See AMT templates in Appendix G.2.
For each domain, we report aggregated and template-wise performance of InstructGPT, ChatGPT and GPT4 as well as human baseline performance in Table 4. Random guessing performance is 54% instead of 50% because of the 13 probing templates.
@@
The performance drop of GPT4 is as high as 4% in zero-shot setting and as high as 5% in COT setting. Moreover, the long-tail performance of models have a marked gap compared to human baselines, being as high as 19% in the “Natural Properties” domain.
The positive template performance experience a “long-tail rise” for all models and most domains.
For all models, the performance on positive templates is significantly lower than on negative templates. Besides “Natural Properties” domain, we witness an increase in performance in the long-tail distribution on the positive templates, contrary to our observation on the negative templates. We examine the rationale the models generate during COT in Table 12 and find that the models tend to avoid drawing a “definite conclusion”. We hypothesize that this phenomenon may be a result of model adjustments made with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), and these model adjustments may have a weaker effect on the long-tail distribution due to the knowledge being less frequent during training. This hypothesis can be further supported by the significant increase in positive template performance after COT prompting, where COT weakens the model’s original distribution over the answer tokens in the in-context examples. We leave the exact explanation to future work.
@@
Performance drop in the long-tail distribution does not occur to humans for 3 out of 4 domains. It is expected because humans can verify their knowledge using search engines, so infrequent knowledge does not challenge humans as much as models. The exception with the locational domain may be due to some relations being less available online(eg. banned_in(Food, Country)).
@@ 5 Related Work
The community has begun to realize the importance of testing language models’ abilities in the long-tail distribution (Godbole and Jia, 2022). Recent works reveal that the performance of LLMs is affected by input data probability. (McCoy et al., 2023; Razeghi et al., 2022), and more works have focused on generating less common data for probing LLMs. RICA (Zhou et al., 2020) proposes to include novel entities in self-contained commonsense statements to evaluate robust inference capabilities. UnCommonSense (Arnaout et al., 2022) proposes to evaluate models on informative negative knowledge about everyday concepts in addition to positively expressed commonsense knowledge. Razeghi et al. (2022) observe a correlation between the model performance on math problems and the frequency of numeric and temporal terms from those instances in the pretraining data.
In addition to probing models’ reasoning ability on less common data, recent works also propose to test language models’ ability to generate less common dataChen et al. (2023) propose a negative knowledge generation task where models generate uncommon knowledge with negation conditioned on constrained keywords. Tang et al. (2023) introduce the “less likely brainstorming” task that asks a model to generate outputs that humans think are relevant but less likely to happen.
Generating uncommon data is not only challenging for language models, but also challenging for humans because of our cognitive biasTversky and Kahneman (1974) observe that humans are prone to more systematic errors when facing uncertain events, and Tversky and Kahneman (1973) reveal that humans tend to evaluate the frequency of classes or the probability of events by availability, i.e., by the ease with which relevant instances come to mind. These traits make it difficult for humans to come up with novel associations (Kray et al., 2006), a crucial ability to create data in the long-tail distribution.
6 Conclusion and Future Work
This paper is situated in the problem of generating challenging evaluation data for LLMs in the long-tail distribution. We proposed a long-tail data generation framework, LINK, which grounds the statements in symbolic rule templates. We showed that zero-shot long-tail knowledge generation is still hard for the SOTA LLMs, and our approach LINK is effective for reaching the long-tail distribution while preserving factual correctness. We also showed that LLM’s reasoning capabilities are especially affected in the long-tail regime.
Using LINK, we created a dataset, LINT, which contains 417 symbolic rules and 54K long-tail knowledge statements. LINT is a useful resource for curating challenging evaluation data for LLMs. Using a simple entailment classification task, we showed that the long-tail data reveals more flaws of LLM’s reasoning capabilities.
LINK can facilitate future work on data augmentation. One can take natural language statements from existing benchmarks (or new benchmark proposals) and turn them into symbolic rules, from which we can generate long-tail data using LINK. Given such long-tail data are less frequently seen during training, it is also useful for mitigate benchmark leakage issues during evaluation.

### v3

@@ 3.1 Instruction-Only Knowledge Generation
We generate knowledge statements from a subset of 200 symbolic rules from LINT using ChatGPT and GPT4 by only providing it with an instruction. We prompt the LLMs with a natural language version of the symbolic rule with data types of the variables specified, and ask them to populate the rule with all variables simultaneously(Table 2). Generations from this prompt serve as the head distribution baseline, to which we compare model generations when they are instructed to generate in the long-tail distribution.
To instruct models to generate long-tail knowledge from symbolic rules, we append “Use less frequent terms of A and B and C” in the prompt. [^5] We tried 10 prompts as shown in Appendix F and they have similar effect on model behavior.
For each rule, we obtain 200 statements using default instruction and 200 statements using long-tail instruction, from ChatGPT and GPT4 respectively. We present our findings below.
3.2 LINK Generations Consistently Fall in Long-tail Distribution
Following Godbole and Jia (2022)’s definition of long-tail statements, we use the most capable LLM that was producing log likelihood at the time of experiments (text-davinci-003) to assign likelihood to generated data. We compare how different models assign distributions to the same statements, in order to ensure that the distribution stays consistent among all models despite the absolute log likelihood difference (Appendix E.2).
We calculate the log likelihood over InstructGPT of all statements generated by LINK, compared with instruction-only ChatGPT and GPT4. We calculate $\delta=mean(D(H))-mean(D(L))$ for each set of statements generated from each symbolic rule, where $D(\cdot)$ means the log likelihood distribution of the probability model, $H$ is the set of statements as head distribution baseline, and $L$ is the set of the statements intended to be in the long-tail distribution. For a model to successfully generate long-tail knowledge statements, the “long-tail” set of sentences should be assigned distinguishably lower probabilities than the sentence in the head distribution.
@@ 3.2 LINK Generations Consistently Fall in Long-tail Distribution
Figure 3 illustrates that while ChatGPT and GPT4 are not able to generate long-tail statements merely from prompting, LINK is able to generate long-tail statements with much lower likelihood. Each grid on the x-axis represents a unique symbolic rule, and each grid on the y-axis represents $\delta\in[-1.5,2]$. A $\delta$ close to 0 means that the intended long-tail distribution generations have the same log likelihoods as the statements in the head distribution, being larger than 0.3 empirically means a decent drop in likelihood, and being negative means the intended long-tail distribution data have even high likelihood than the head distribution data.
Averaged across 200 sampled rules, LINK has a positive $\delta$ of 0.48, while ChatGPT and GPT4 each has a delta of -0.14 and -0.02. The $\delta$ values for LINK (blue line) float above 0 for most of the rules, some even being above 0.5. On the other hand, $\delta$ values for ChatGPT (red line) and GPT4 (green line) mostly locate around 0, with many being negative.
To better illustrate the distribution of statements generated by LINK, ChatGPT and GPT4, we plot the log likelihood of the generated statements for one symbolic rule from the three methods Figure 4. To eliminate noise from incorrect statements on the distribution, we only plot the statements that are marked as correct in human evaluation (explained in § 3.3). The likelihood distribution of more rules can be found in Figure 9.
The long-tail statements from LINK clearly fall in a much lower probability distribution than GPT4’s “long-tail” generations. Moreover, GPT4’s generation in the “long-tail distribution” in fact falls in the same probability distribution as its head distribution generations.
3.3 LINK Achieves Higher Data Correctness than Instruction-Only LLMs
@@
In addition to distribution correctness, we also evaluate data type conformity and factual correctness of LLMs’ long-tail knowledge generations using crowdworkers from Amazon Mechanic Turk (AMT). For data type conformity, we ask an AMT worker Is {variable} a {data type}? for each variable in the symbolic rule. For factual correctness, we ask an AMT worker Does the premise entail the conclusion? We sample 4,000 statements from LINT for human evaluation, of which 2,025 are from head distribution and 1,975 are from long-tail distribution. We take 3 annotations for each statement and take the majority vote. Annotator agreement can be found in Appendix I.3. The AMT template can be found in Appendix I.2.
Table 3 shows that instruction-only ChatGPT and GPT4 underperform LINK in both data type conformity and factual correctness. Without LINK, both models struggle more with factual correctness, a foreseeable behavior in the low likelihood realm. For domain wise performance see Table 17. For examples of failure cases, see Table 20.
3.4 Ablation Studies
Our results above show that LINK is better than instruct-only generations with ChatGPT and GPT4, in terms of both “long-tailedness” and factual correctness. In this section, we perform ablation studies on the role of reranker, critic, and knowledge models in LINK.
Ablation on reranker.
Our variable-wise reranker is essential for pushing LINK generations into the long-tail distribution. Figure 6 presents the distribution comparison of generated statements by LINK and the variant without the reranker. Without the reranker, the statements for the long-tail distribution are pulled towards the head distribution, making the two completely inseparable.
Post-hoc reranking over LLM generations does not have the same effect as variable-wise reranking in LINK. Figure 7 illustrates the distribution of generated statements by LINK, compared to instruction-only GPT4 and instruction-only GPT4 reranked by InstructGPT. Post-hoc reranking barely changes the distribution of generations, even when using the same model as the evaluation.
Even though log likelihood is both used by reranker and evaluation, they are taken from different models and using different inputs. The statements we use for ranking the knowledge beams are shorter than the final statement, as they only consist of partial predicates. Despite these differences, variable-wise reranking that uses a smaller model achieves the separation that post-hoc filtering with the evaluation model cannot achieve.
Our findings highlight the importance of performing variable-wise reranking in LINK.
@@ 3.4 Ablation Studies
Critic models are essential for guaranteeing the generation quality, especially in the long-tail distribution. Table 8 in Appendix D.2 shows that removing the critic and removing both reranker and critic leads to significant drops of data type conformity and factual correctness of generations in the long-tail distribution. Note that LINK w/o reranker + critic has higher generation quality than LINK w/o critic. This is because without a reranker the model is only able to generate statements in the head distribution, making it easier to be factually correct than in the long-tail distribution. This observation further suggests that critic models are essential for generation qualities in the long-tail distribution.
Ablation on Knowledge Model.
Our analysis above has shown that by simply providing an instruction to ChatGPT and GPT4, we cannot effectively generate knowledge statements that are both high quality and in the long-tail distribution. By adding LINK, we can improve both distribution correctness and generation quality, as shown in Figure 8 and Table 10.
@@
Domain Distribution Llama2-70B ChatGPT GPT4 Human Baseline Pos Neg All Pos Neg All Pos Neg All All Natural Properties Head 2.78 6.72 0.44 1.45 5.56 0.0 13.68 43.83 9.23 82.31 Long-tail 2.64 3.49 0.0 1.08 5.29 0.0 10.82 39.9 7.21 82.45 $\Delta$100% - - -0.0% - - -21.89% 0.17% Temporal Head 6.58 2.6 0.46 10.72 38.28 4.13 68.3 43.95 36.91 84.69 Long-tail 7.43 3.07 0.0 9.05 32.15 2.26 60.58 38.93 28.11 83.20 $\Delta$100% - - -45.28% - - -23.84% -1.76% Outcomes and Effects Head 7.62 7.93 1.22 19.92 29.27 6.1 57.32 55.18 41.16 83.83 Long-tail 8.5 9.97 0.59 19.65 26.1 2.64 55.72 46.63 33.43 85.13 $\Delta$51.64% - - -56.72% - - -18.78% 1.56% Locational Head 8.57 7.14 0.0 17.14 10.0 5.71 18.57 18.57 2.86 75.71 Long-tail 11.24 8.99 3.37 15.73 4.49 1.12 37.08 8.99 3.37 67.42 $\Delta$ 0.0% - - -80.39% - - 17.83% -10.95% Total Head 5.08 5.28 0.56 8.21 20.67 2.62 39.49 44.87 23.64 83.12 Long-tail 5.69 4.67 0.27 7.76 17.86 1.28 36.58 39.34 18.66 82.44 $\Delta$51.79% - - -51.15% - - -21.07% -0.82%
@@
Table 4: Performance on the entailment classification task of three LLMs decreases on the long-tail distribution compared to the head distribution, while human performance does not.
4 LLMs’ (Lack of) Generalization in the Long-tail Distribution
Using data from LINT, we evaluate LLM generalization through an entailment classification task on inferential knowledge in the long-tail distribution.
@@ 4 LLMs’ (Lack of) Generalization in the Long-tail Distribution
We use all human evaluated knowledge statements except for those with incorrect data types in LINT, with 1,925 statements in head distribution and 1,856 statements in long-tail distribution. Statements rated as factually correct has entailment between premise and conclusion, and statements rated as factually incorrect has contradiction between premise and conclusion.
@@
In order to prevent LLMs’ template biases from misleading the evaluation, we convert each statement into 13 question templates, where each question templates corresponds to a positive label (“Yes”, “True”, or “Right”) or a negative label (“No”, “False”, or “Wrong”). The question templates are summarized in Appendix G.1. We consider a model answering accurately about one statement only if the model answers all question templates correctly.
We evaluate three LLMs: llama2-70B, ChatGPT and GPT4. In order to enforce the model to predict the target token sets and minimize format noncompliance, we use Chain-of-Thought (CoT) (Wei et al., 2022) prompting that includes 2 in-context examples with randomly shuffled orders of positive label and negative label.
For each domain, we report aggregated (All) performance of each model as well as human baseline performance in Table 4. We also include performance on positive labels only and negative labels only. We also mark relative performance drop $\delta=\frac{t-h}{h}$, where $h$ and $t$ are head and tail distribution aggregated performance.
@@
We obtain human performance on the same set of statements. We recruit 17 AMT workers who do not participate in the evaluation task (and thus have not seen the task data). The workers see the knowledge statements in premise, conclusion format and are asked to select “yes/no” to whether the premise entails the conclusion. The workers are asked to use search engines to verify their answers. See AMT templates in Appendix I.2.
We make the following observations on LLM generalization in long-tail NLI.
@@
All models exhibit a large relative drop in performance in the long-tail distribution. The most competitive model, GPT4, has a 21% overall drop from head to long-tail distribution, while other models exhibit a even larger drop.
@@
Performance drop in the long-tail distribution does not occur to humans for 3 out of 4 domains. It is expected because humans can verify their knowledge using search engines, so infrequent knowledge does not challenge humans as much as models (discussion in Appendix A). The exception with the locational domain may be due to some relations being less available online(eg. banned_in(Food, Country)).
Brittleness towards question templates.
The huge gap between model and human baseline performance indicate that LLMs cannot reliably reason on the same statement when question templates change. We find that model performance between positive and negative labels can be very different for certain domains, indicating that models are miscalibrated for positive and negative answer tokens. Although this phenomenon is not entirely due to the shift of distribution of the knowledge statements, it is caused by model’s unfamiliarity of certain question templates. For example, we find that template 5 and template 12 are the same question (“Premise: … Conclusion: … Does the Premise entail the Conclusion?") with opposite answers (“Yes" and “No"), but all models’ performance on template 5 is significantly lower than that on template 12. This suggests that models are more likely to create false negatives in such context, another evidence of performance drop due to long-tail distribution.
In summary, our analysis shows that while human’s inferential reasoning is not affected by their familiarity to the data (provided that they know the entities involved), model’s inferential reasoning ability drops over long-tail knowledge. Our result highlights the importance and effectiveness of long-tail evaluation for model generalization, and our dataset LINT can be used as a useful resource for testing inferential generalization of LLMs.
@@ 5 Related Work
Works on model generalization analysis have focused on generating adversarial examples for model evaluation (Zhang and Li, 2019; Ziegler et al., 2022; Perez et al., 2022; Casper et al., 2023), flagging abnormal inputs that are likely to trigger bad behavior. Recently, the community has realized the importance of testing language models’ abilities in the long-tail distribution (Godbole and Jia, 2022). Works reveal that LLM performance is affected by input data probability. (McCoy et al., 2023; Razeghi et al., 2022), and more works have focused on generating less common data for probing LLMs. RICA (Zhou et al., 2020) proposes to include novel entities in self-contained commonsense statements to evaluate robust inference capabilities. UnCommonSense (Arnaout et al., 2022) proposes to evaluate models on informative negative knowledge about everyday concepts in addition to positively expressed commonsense knowledge. Razeghi et al. (2022) observe a correlation between the model performance on math problems and the frequency of numeric and temporal terms from those instances in the pretraining data.
In addition to probing models on less common data, recent works also test LLMs generating less common dataChen et al. (2023) propose a negative knowledge generation task where models generate uncommon knowledge with negation conditioned on constrained keywords. Tang et al. (2023) introduce the “less likely brainstorming” task that asks a model to generate outputs that humans think are relevant but less likely to happen.
Generating uncommon data is challenging not only for LLMs, but also for humans because of our cognitive biasTversky and Kahneman (1974) observe that humans are prone to more systematic errors when facing uncertain events, and Tversky and Kahneman (1973) reveal that humans tend to evaluate the frequency of classes or the probability of events by availability, i.e., by the ease with which relevant instances come to mind. These traits make it difficult for humans to come up with novel associations (Kray et al., 2006), a crucial ability to create data in the long-tail distribution.
6 Conclusion
Using NLI as a case study, we illustrated the significant potential of long-tail data in uncovering the generalization limitations of LLMs. We introduced the first systematic framework designed to generate inferential data within the long-tail distribution, and then demonstrated a noteworthy performance drop of LLMs in the long-tail examples. Our work initiates a new line of research focused on long-tail data discovery and generation, urging the research community to adopt long-tail evaluation in the development of LLMs.